OpenStudy (anonymous):

please help me solve (6x^13)^3

OpenStudy (anonymous):

216x^39

OpenStudy (anonymous):

You find 6^3 = 6 * 6 * 6 = 216 When you take a power to a power you multiply so x^13 to the 3rd power is x^39 because 13 * 3 = 39 so 216 x^39